OpenStudy (anonymous):

Medal + Fan There were 20 marbles in a bag, of which 15 were blue, and the rest were white. For a game, 35 of the 20 marbles were chosen. Use this information to answer the questions below. If not enough information is given to answer a question, click on "Not enough information."

OpenStudy (anonymous):

How many of the bag's blue marbles were not chosen?

OpenStudy (anonymous):

how many of the bag marble were chosen?

OpenStudy (anonymous):

@tester97 @cheerleader @chucky00 @kcurri @protected @amandanewsomm @Ruu @fxck.hxnnah @vlltimelow @kym02 @Destinymasha @damien @cj49 @XAmberX @zepdrix @TheSmartOne

OpenStudy (anonymous):

That doesn't even make sense. How will you get 35 from 20? ._.

OpenStudy (anonymous):

it 3/5

OpenStudy (anonymous):

@Ruu

OpenStudy (anonymous):

Is that what you got?

OpenStudy (anonymous):

no the 35 is 3/5

OpenStudy (nightwielder):

3/5 * 20=12 12 marbles are drawn. At the most, 12 blues can be drawn. If that is the most then having 3 marbles not drawn is the least. Also, if all 5 whites were drawn then 12-5=7 at least blue marbles taken. Meaning blues not taken at most is 8. Answer: \[3\le x \le8\]